the incidence of occupational disease in an industry is such that the workman has 20 percentage of suffering from it. what is the probability of that out of five workman selected ?
1) 2 will contact the disease
2) none will suffer
3)3 or more will contact the disease

Answer by Boreal(15235)

5C2 or 10 are the number of ways 2 can be selected.
multiply by 0.2^2*0.8^3=0.2048
------------
none is 0.8^5=0.3277
-------------------
3 or more will contact
1 will is 5*0.2*0.8^4=0.4096 (the expected value)
0,1,2 add to 0.93=421, so 3,4,5 are 0.0579.
Can check by
3: 10*0.2^3*0.8^2=0.0512
4:5*0.2^4*0.8=0.0064
5:0.2^5=0.0003
add to 0.0579